OpenStudy (anonymous):

Solve the first order equation y_(k+1)=ky_k+k^2 recursively, given the initial condition y_0=1.

OpenStudy (anonymous):

\[y _{k+1}=ky _{k}+k ^{2}\]

OpenStudy (anonymous):

initial condition \[y _{0}=1\]

OpenStudy (anonymous):

hi friend, do you have any idea?

OpenStudy (anonymous):

i try to use nonhomogeneous recurrence relation method...but i didnt get the right answer...

OpenStudy (anonymous):

\[y _{k}=ky _{k-1}+k ^{2}\] solve homo part: \[y _{k}=ky _{k-1}\] we get \[r ^{1}=ky _{0}=k\] \[r=k , r-k=0\] thus,\[y _{k}^{h}=c _{1}k ^{k}\]

OpenStudy (anonymous):

for nonhomo \[F(n)=k ^{2}\] so F(n) in the form of equation \[P(k)=ak ^{2}+bk+c\]

OpenStudy (anonymous):

so we solve by letting \[P(k)=y _{k}\] compare both side, then find value a, b and c

OpenStudy (anonymous):

the i get the solution for non homo part \[F(n)=\frac{ 2 }{ 5} k ^{2}+\frac{1}{ 5 }k\]

OpenStudy (anonymous):

after combining and find \[c ^{1}\] still dont get correct answer...

OpenStudy (anonymous):

hold on. I think the LHS has problem. the characteristic equation you get from the original one may be wrong.

OpenStudy (anonymous):

I got L^2 -kL = 0 (homogenous part) ---> L (L -k) = 0 ---> L =0 and L =k. Ok, sorry, end up at the same result, just different way. ok, go ahead

OpenStudy (anonymous):

I assume that yours is right, so the combination of both homogenous and nonhomogenous part is?

OpenStudy (anonymous):

y\[y _{k}=c _{1}k ^{k}+\frac{ 2 }{ 5 }k ^{2}+\frac{ 1 }{5 }k\] find c1 by subtituting \[y _{1}=0\] we get c1=-3/5

OpenStudy (anonymous):

your initial condition is yo not y1?

OpenStudy (anonymous):

yup...but you can find y1 by substitute k=0 in the original eq...

OpenStudy (anonymous):

sory let k=1 and you get y2=0

OpenStudy (anonymous):

y2=1 sorry...

OpenStudy (anonymous):

no, not that, you don't have C yet, how can you claim the form and then using your claim to calculate something else?

OpenStudy (anonymous):

My prof said that we have to calculate everything based on the given condition, claim the equation and use it to check for every values of k

OpenStudy (anonymous):

from the original eq \[y _{k+1}=ky _{k}+k ^{2}\] so i substitute k=1 then get y2=1

OpenStudy (anonymous):

i find the value of y starting with k=0 then k=1 and so on...

OpenStudy (anonymous):

don't forget yo means k = -1, should we step back one term? to get the series starts at k =0?

OpenStudy (anonymous):

hurm...

OpenStudy (anonymous):

Maybe I am wrong, but your k doesn't match with the y at all. if this way wrong, try another way. one more question, did you try generating function for this stuff?

OpenStudy (anonymous):

It's never never simple, but what we study, try again friend, don't give up. we still have our profs, right?

OpenStudy (anonymous):

when do you need this stuff? I can ask my prof via email. I will send you the answer right after I get it, is it ok?

OpenStudy (anonymous):

this is my assingment...so my duedate is tommorow 4 pm....my lecturer just give us 1 day to finish it....

OpenStudy (anonymous):

but anyway....thanks for ur support..... =) i will try finding another way....

OpenStudy (anonymous):

hey.. this sounds straight forward!!

OpenStudy (anonymous):

start with k = 0; \(y_1=(0)y_0+(0)^2=0\) k=1 \(y_2=(1)y_1+(1)^2=1\) k=2 \(y_3=(2)y_2+(2)^2=6\) k=3 \(y_4=(3)y_3+(3)^2=27\) and so on

OpenStudy (anonymous):

the given equation is itself the solution. it is not given as a differential equation.

OpenStudy (anonymous):

yes....but we need to find the solution in the form of y(k) or something...but not in recursive form...

OpenStudy (anonymous):

so, "y" is a function of "k"??

OpenStudy (anonymous):

i'm not sure wether k is a constant or an unknown...thats make me confuse....

OpenStudy (anonymous):

should be \[y _{k+1} - ky _{k}=k ^{2}\]

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

ok....then....

OpenStudy (anonymous):

\[ y_{k+1}=ky_k+k^2\\\;\\ y_{k+1}=k[(k-1)y_{k-1}+(k-1)^2]+k^2=k(k-1)y_{k-1}+k^2+k\times (k-1)^2\\ \implies y_{k+1}=k(k-1)y_{k-1}+k[k+(k-1)^2]\\\;\\ y_{k+1}=k(k-1)[(k-2)y_{k-2}+(k-2)^2]+k[k+(k-1)^2]\\ \implies y_{k+1}=k(k-1)(k-2)y_{k-2}+k(k-1)(k-2)^2+k[k+(k-1)^2]\\ \implies y_{k+1}=k(k-1)(k-2)y_{k-2}+k[k+(k-1)[(k-1)+(k-2)^2]\\;\\ y_{k+1}=k!\times y_1+k\sum_{n=1}^kn\\\;\\ y_{k+1}=k!\times y_1+k\frac{k(k+1)}{2} \]

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

This is a question for the smart people.

OpenStudy (anonymous):

did I make sense in the last one?

OpenStudy (anonymous):

@Mertsj could you bring in the task force? :)

OpenStudy (mertsj):

Try Phi or Satellite or Turing Test

OpenStudy (anonymous):

@phi @ParthKohli calling all the backup!! heavy artillary

OpenStudy (anonymous):

@hartnn @Callisto @Luis_Rivera Calling chopper backup here

OpenStudy (mertsj):

@amistre64 @satellite73

OpenStudy (anonymous):

@Mertsj they both offline!

OpenStudy (mertsj):

But one never knows when they will show up and see their notifications.

OpenStudy (mertsj):

@TuringTest

OpenStudy (anonymous):

My base willl be destroyed by then!

OpenStudy (anonymous):

@narsha what if we take the "Z" transform of this thing? it'd be way easier that way!

OpenStudy (anonymous):

"Z" transform...meaning????

OpenStudy (anonymous):

transform the system to complex plane.

OpenStudy (anonymous):

??? kind of confusing....now im trying to use method of generating function because i stuck with nonhomogeneous recurrence relation.....

OpenStudy (anonymous):

is there anyway to check the answer?

OpenStudy (anonymous):

@amistre64 could you look at this pls

OpenStudy (anonymous):

@electrokid , can i ask something about the answer that u just type above....how u generate from \[y _{k-2}\] to \[y _{1}\] ?????????

OpenStudy (anonymous):

my bad. I did not use it. and deleted that post.

OpenStudy (anonymous):

my logic starts right after that!

OpenStudy (anonymous):

i think when u use back substitution is correct....but i dont get how u simplify from \[y _{k+1}=k(k−1)(k−2)yk−2+k[k+(k−1)[(k−1)+(k−2)2]\] becomes \[y _{k+1}=k! \times y _{1}+k(\frac{ k(k+1) }{ 2 })\]

OpenStudy (anonymous):

left part, as you keep substituting, you reach the k!

OpenStudy (anonymous):

the second part, I am not sure. I thought it'd reach the artimatic series sum.

OpenStudy (anonymous):

\[ y_{k+1}=k(k-1)(k-2)[(k-3)y_{k-3}+(k-3)^2]+k[k+(k-1)[(k-1)+(k-2)^2]\\ \quad=k(k-1)(k-2)(k-3)y_{k-3}+k[k+(k-1)[(k-1)+(k-2)^2+(k-2)(k-3)^2]\\ \quad=k(k-1)(k-2)(k-3)y_{k-3}+k[k+(k-1)[(k-1)+(k-2)[(k-2)+(k-3)^2] \]

OpenStudy (anonymous):

how can the second part be represented!!!!

OpenStudy (anonymous):

@narsha do you see a pattern there?

OpenStudy (anonymous):

\[\Large{ y_{k+1}=k(k-1)(k-2)(k-3)y_{k-3}+\\\qquad\quad\color{red}{k[k+(k-1)[(k-1)+(k-2)[(k-2)+(k-3)^2}}] \]

OpenStudy (anonymous):

\[ k^2+k(k-1)^2+k(k-1)(k-2)^2+k(k-1)(k-2)(k-3)^2+\dots \]

OpenStudy (anonymous):

i see that pattern....but at thelast answer, how you get \[y _{1}\] ???

OpenStudy (anonymous):

ooooh \[ =\sum_{n=1}^k (k-n)!(n)^2 \]

OpenStudy (anonymous):

\[ y_{k+1}=k!\times y_1+\sum_{n=1}^k (k-n)!(n)^2 \]

OpenStudy (anonymous):

noo. has to be rational.

OpenStudy (anonymous):

\[ =\sum_{n=1}^k {k!\over n!}(n)^2 \]

OpenStudy (anonymous):

yep. this satisfies good.

OpenStudy (anonymous):

\[\boxed{\Large y_{k+1}=k!\times y_1+\sum\limits_{n=1}^k \huge \frac{k!\times n^2}{n!} }\]

OpenStudy (anonymous):

I can do back substitution till y_1. what happens when we do the next one to get y_0?

OpenStudy (anonymous):

ermmm....\[y _{0}\] is already given as initial cond....so i dont think we need to find it....just go on with y_2 , y_3 and so on....

OpenStudy (anonymous):

nono. the general solution \(y_{k+1}\) should be explressed in terms of \(y_0\)

OpenStudy (anonymous):

if we want to express \[y _{0}\] then we have to let k=-1.....but here k should start with 0.. cause when we substitute k in original eq

OpenStudy (anonymous):

\[y _{k+1}=ky _{k}+k ^{2}\] we cant find it

OpenStudy (amistre64):

\[y _{k+1}=ky _{k}+k ^{2}~:~y_0=1\] \[y _{1}=0y _{0}+0 ^{2}\] \[y _{2}=1(y _{0})+1^2\] \[y _{3}=2(y _{0}+1^2)+2^2\] \[y _{4}=3(2(y _{0}+1^2)+2^2)+3^2\] \[y _{5}=4(3(2(y _{0}+1^2)+2^2)+3^2)+4^2 \] \[y _{6}=5(4(3(2(y _{0}+1^2)+2^2)+3^2)+4^2)+5^2\\~~~~=5\cdot4\cdot3\cdot2\cdot y_0+5\cdot4\cdot3\cdot2\cdot 1^2+5\cdot4\cdot3\cdot2^2+5\cdot4\cdot3^2+5\cdot4^2+5^2 \] \[y _{n}=(n-1)!\cdot y_0+(\frac{1(n-1)!}{0!} +\frac{2(n-1)!}{1!}+\frac{3(n-1)!}{2!}+...+\frac{(n-1)(n-1)!}{(n-2)!}) \] \[y _{n}=(n-1)!~y_0+(n-1)!~(\frac{1}{0!} +\frac{2}{1!}+\frac{3}{2!}+...+\frac{(n-1)}{(n-2)!})~:~y_0=1\] \[y _{n+1}=4~n!+n!~(\frac{3}{2!}+\frac{4}{3!}+...+\frac{n}{(n-1)!})\] @electrokid if you havent gotten it by now, then i think your on the right track. Im sure i missed something while trying to keep track of the latex